OpenStudy (anonymous):

I'm having trouble understanding how to differentiate sinxtanx. Can someone please help me?

OpenStudy (baldymcgee6):

For this we need the product rule

OpenStudy (anonymous):

Im stuck after secxsec(^2)x+tanx(secx+tanx)

OpenStudy (baldymcgee6):

what is the derivative of sin(x)?

OpenStudy (anonymous):

cosx

OpenStudy (baldymcgee6):

and the derivative of tan(x)?

OpenStudy (baldymcgee6):

right, so if y= sin(x)*tan(x), then y' = (sin(x))'(tan(x)) + sin(x)(tan(x))'

OpenStudy (anonymous):

\[\sec ^{2}x\]

OpenStudy (baldymcgee6):

good, so now that you know how to differentiate both terms, you can do this y' = (sin(x))'(tan(x)) + sin(x)(tan(x))' = ???

OpenStudy (anonymous):

I'm not sure how my solutions manual got secx(1+2tan(^2)x) though

OpenStudy (anonymous):

OHH wait. I got it. Thank you! I overlooked a couple steps.

OpenStudy (baldymcgee6):

good job, please leave me a medal and fan :D